Simplify these expressions and find their values if x = 3, a = – 1, b = – 2.
 2a – 2b – 4 – 5 + a

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that a = -1, b = -2
We have,
= 2a + a–2b–4 - 5
= 3a–2b - 9
Then, substitute the value of a and b in the equation
= (3 × (-1))–(2 × (-2))–9
= -3–(-4)–9
= - 3 +4–9
= -12 + 4
= -8

Thanks